OpenStudy (anonymous):

Solve by the elimination method.... I am confused x+y=-5 6x+y=-10 5x+5y=-11 7x-2y=17 7x-6y=-2 6x+7y=59 4x+5y=-7 -4x+y=13

OpenStudy (lgbasallote):

...you confuse me...

OpenStudy (anonymous):

sorry , i am trying to solve each problem using the elimination method

OpenStudy (callisto):

Is it one question or more than one question there?

OpenStudy (anonymous):

Elimination means you multiply one equation by a number and add that to the other equation in the pair to try and cancel out one of the variables. Assuming the first two equations are paired, you can multiply the first equation by -1 to get: -x-y=5 Then add it to the second equation: -x+6x-y+y=5+10 which simplifies to 5x=15,so x=3. Now you can plug in 3 for x in either equation, so I'll use the first: 3+y=-5, so y=-8. The other pairs of equations would be similar, but you would have to multiply equations by different numbers, and sometimes you have to multiply both equations by two different numbers to eliminate a variable.